OpenStudy (anonymous):

A mass is oscillating on a spring with a period of 1.20 s. At t = 0 the mass has zero speed and is at x = 4.85 cm. What is the magnitude of the acceleration at t = 3.30 s?

OpenStudy (mos1635):

V=0 =>X=Xmax=maggnute=A=4.85^-2 m T=1.2 sec ω=2π/Τ α=-ω^2 * Α sin(ω t + π/2) go for it

OpenStudy (anonymous):

does this stand for the acceleration am trying to find α?

OpenStudy (anonymous):

am getting (-4.780cm/s^2) though not sure,i have one trial left

OpenStudy (anonymous):

u there?

OpenStudy (mos1635):

yes

OpenStudy (anonymous):

is this α my acceleration?

OpenStudy (anonymous):

i too \[w=2\pi f\]

OpenStudy (anonymous):

and A=4.85^-2 m

OpenStudy (anonymous):

α=-ω^2 * Α sin(ω t + π/2) this is my amplitude A=4.85^-2 m ? and this is my \[w=2\pi f\] and my acceleration is this α?

OpenStudy (mos1635):

yes

OpenStudy (mos1635):

a=4.781906 m!!!!/sec^2 that is my ans

OpenStudy (anonymous):

meaning i should x=4.85cm to meters?

OpenStudy (mos1635):

I am used to SI

OpenStudy (anonymous):

ok so the acceleration uis the negative of the answer u just wrote,-4.781?

OpenStudy (mos1635):

i suppose if given x to cm acceleretion will be acceptable to -478.1906 cm/sec^2 (yes minus)

OpenStudy (anonymous):

oops! the system says it was wrong,and dat was my last trial :)

OpenStudy (mos1635):

you did nt put the whole value? -478.20 should be our ans according to given values i wrote the accurate number as i am used to (sorry)